OpenStudy (anonymous):

find the derivative of: f(t)=9tan(t)+7csc(t) my answer is: 9sin3x-7cos2x all over cosxsin3x just wondering if thats right.

OpenStudy (anonymous):

[ 9tan(t) ]' = 9/cos^2(t) [7ccs(t)]' = -7cos(t)/sin^2(t)

OpenStudy (anonymous):

9⋅(sec(t))^2−7⋅cot(t)⋅csc(t)

OpenStudy (anonymous):

yes same as i wrote

OpenStudy (anonymous):

so the only thing i have wrong is that i need to put cos2 in the denominator?

OpenStudy (anonymous):

sin3x is sin^3(x) for you ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it still doesnt make it right since the second term will be missing stuff

OpenStudy (anonymous):

well for my derivative i have 9sec^2(x)+7(-csc2x)cot^2(x)

OpenStudy (anonymous):

9/cos^2(t) - -7cos(t)/sin^2(t) = [9sin^2(t) - 7cos^3(t)] / sin^2(t)cos^2(t)

OpenStudy (anonymous):

see i had sin^3 in my denominator

OpenStudy (anonymous):

so there is something wrong with the second term since : [7ccs(t)]' = -7cos(t)/sin^2(t) or we can represent it as -7cot(t)csc(t)

OpenStudy (anonymous):

yes that's what i have, then i changed it all into sin and cos.

OpenStudy (anonymous):

but you wrote 7(-csc2x)cot^2(x)

OpenStudy (anonymous):

this is not the same

OpenStudy (anonymous):

i dont know why i wrot that hah.

OpenStudy (anonymous):

lol so show me your answer before manipulation

OpenStudy (anonymous):

now my answer is: \[\frac{ 9\sin ^{2}x-7\cos ^{2}x }{ \cos ^{2}xsin ^{2}x }\]

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

you should have 9sin^2(x) -7cos^3(x) upstairs

OpenStudy (anonymous):

but the bottom is correct? i see what i did wrong on the top now when i got my lcd i didnt add the other cos x to the 2cos^2(x)

OpenStudy (anonymous):

the bottom correct but the numerator isnt

OpenStudy (anonymous):

so my numerator should be 9sin^2(x) -7cos^3(x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and it should be t instead of x ;)

OpenStudy (anonymous):

so done :) ?